OpenStudy (anonymous):

Simplify: 20xy(x+y)=

OpenStudy (anonymous):

\[20(x ^{2}y+xy^{2})\]

OpenStudy (anonymous):

just distribute

OpenStudy (anonymous):

20x^2y +20xy^2

OpenStudy (anonymous):

How come not 20x^2y^2?

OpenStudy (anonymous):

Got it. Thank you.

OpenStudy (unklerhaukus):

\[20xy(x+y)=(20xy\times x)+(20xy\times y)=20x^2y+20xy^2\]

OpenStudy (anonymous):

???Should 20 be in the solution twice?

OpenStudy (anonymous):

Ok